OpenStudy (anonymous):

Which of the following is an example of a parent function? A.y= x^3+x^2+2x B.y = |x| C. y=2/x D. y=3x^2

OpenStudy (ranga):

A parent function is a set of basic functions used as building blocks for more complicated functions. Examples include: y = x, y = 1/x, y = x^2, y = |x|, etc. Here it is B. y = |x|

OpenStudy (anonymous):

ok! thank you!

OpenStudy (ranga):

you are welcome.